OpenStudy (anonymous):

what's the answer to 9x+18y=0 step by step

OpenStudy (anonymous):

solving for both x , y?

OpenStudy (anonymous):

in slope intercept form

OpenStudy (dumbcow):

solve for y subtract 9x on both sides 9x +18y =0 -9x -9x ---------------- 0 +18y = -9x divide by 18 on both sides 18y/18 = -9x/18 y = -9x/18 reduce fraction y=-x/2 slope=-1/2 y_intercept = 0

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

is the answer to "7x-1=37 +8" y=7/3x - 3? y intercept form

OpenStudy (dumbcow):

is that 3y +8 ?? if so then you are correct good job